# A lower bound for the Weisfeiler-Leman dimension of circulant graphs

### v1

@@ 3. Binary schemes
Let ${\cal X}$ be a scheme on $\Omega$. For any integer $m\geq 1$ and a tuple $x\in\Omega^{m}$, we define the $m\times m$ array
[…]
The scheme ${\cal X}$ is said to be binary if for any $m$ and any two $m$-tuples $x,y\in\Omega^{m}$ such that $R_{{\cal X},m}(x)=R_{{\cal X},m}(y)$, there exists an automorphism of ${\cal X}$, that takes $x$ to $y$. In this case, the group $\operatorname{Aut}({\cal X})$ is transitive, and acts transitively on each basis relation of ${\cal X}$.
@@
Below, ${\cal X}$ is a scheme on $\Omega$, $m\geq 1$ is an integer, and $x,y\in\Omega^{m}$ are such that $R(x)=R(y)$, where $R=R_{{\cal X},m}$. First, assume that ${\cal X}$ is regular. Then there exists a unique automorphism $f\in\operatorname{Aut}({\cal X})$ such that ${x_{1}}^{f}=y_{1}$. Let $1\leq i\leq m$. By the assumption, the point $y_{i}$ is the unique one for which $r(x_{1},x_{i})=r(y_{1},y_{i})$, where $r=r_{\cal X}$. Since $r(x_{1},x_{i})^{f}=r(y_{1},x_{i}^{f})$, this shows that $y_{i}={x_{i}}^{f}$. Thus $y=x^{f}$ and ${\cal X}$ is binary.
Let ${\cal X}={\cal X}_{1}\otimes{\cal X}_{2}$, where ${\cal X}_{1}$ and ${\cal X}_{2}$ are binary schemes on $\Omega_{1}$ and $\Omega_{2}$, respectively. Denote by $\operatorname{pr}_{i}:\Omega_{1}\times\Omega_{2}\to\Omega_{i}$ the natural projection, $i=1,2$. Identifying in a natural way the sets $(\Omega_{1}\times\Omega_{2})^{m}$ and $\Omega_{1}^{\phantom{\,}m}\times\Omega_{2}^{\phantom{\,}m}$, we can write $x=(\operatorname{pr}_{1}x,\operatorname{pr}_{2}x)$ and $y=(\operatorname{pr}_{1}y,\operatorname{pr}_{2}y)$. By the definition of tensor product, we have
@@
where $R_{1}=R_{{\cal X}_{1},m}$, and $R_{2}=R_{{\cal X}_{2},m}$. Since the scheme ${\cal X}_{i}$ is binary, there exists an automorphism $f_{i}\in\operatorname{Aut}({\cal X}_{i})$ such that $\operatorname{pr}_{i}y=(\operatorname{pr}_{i}x)^{f_{i}}$, $i=1,2$. It follows that the permutation $f:=(f_{1},f_{2})$ of $\Omega_{1}\times\Omega_{2}$ belongs to the group $\operatorname{Aut}({\cal X}_{1})\times\operatorname{Aut}({\cal X}_{2})=\operatorname{Aut}({\cal X})$. Thus,
@@
and the scheme ${\cal X}$ is binary.
Now let ${\cal X}={\cal X}_{1}\wr{\cal X}_{2}$. Then $e={\bf 1}_{\Omega_{1}}\otimes 1_{\Omega_{2}}$ is a parabolic of the scheme ${\cal X}$. Each class of $e$ is of the form $\bar{\alpha}=\Omega_{1}\times\{\alpha\}$, where $\alpha\in\Omega_{2}$. The bijection $\Omega_{2}\to\Omega/e$, $\alpha\mapsto\bar{\alpha}$, $\alpha\in\Omega_{2}$ is an isomorphism from the scheme ${\cal X}_{2}$ to the quotient scheme $\bar{\cal X}={\cal X}_{\Omega/e}$; in particular, the scheme $\bar{\cal X}$ is binary.
@@
where $\bar{R}=R_{\bar{\cal X},m}$, $\bar{x}=(\overline{x_{1}},\ldots,\overline{x_{m}})$, and $\bar{y}=(\overline{y_{1}},\ldots,\overline{y_{m}})$. Thus there exists an automorphism $\bar{f}\in\operatorname{Aut}(\bar{\cal X})$ such that $\bar{y}={\bar{x}}^{\bar{f}}$. Because the natural homomorphism $\operatorname{Aut}({\cal X})\to\operatorname{Aut}(\bar{\cal X})$ is surjective (see [5, Theorem 3.4.6.]), one can find an automorphism $g\in\operatorname{Aut}({\cal X})$ such that $g^{\Omega/e}=\bar{g}$. Replacing $x$ by $x^{g}$ and taking into account that $R(x^{g})=R(x)$, we may assume that
@@
where $R_{\Delta}=R_{{\cal X}_{\Delta},k}$. Because the scheme ${\cal X}_{\Delta}={\cal X}_{1}$ is binary, there exists an automorphism $f_{\Delta}\in\operatorname{Aut}({\cal X}_{\Delta})$ such that $y_{\Delta}=(x_{\Delta})^{f_{\Delta}}$. Let us define a permutation $f\in\operatorname{Sym}(\Omega)$ such that $f^{\Delta}=f_{\Delta}$ for all $\Delta\in\Omega/e$. Then obviously $x^{f}=y$. To complete the proof, it remains to note that $f\in\operatorname{Aut}({\cal X})$, because the kernel of the homomorphism $\operatorname{Aut}({\cal X})\to\operatorname{Aut}(\overline{\cal X})$ coincides with the direct product of $|\Omega/e|$ copies of the group $\operatorname{Aut}({\cal X}_{\Delta})$. ∎
@@ 4.1. Definitions
Let ${\mathscr{R}}$ be a linear ordered set of binary relations on $\Omega$, and let $m\geq 1$ be an integer. The Weisfeiler-Leman algorithm (the $m$-dim $\operatorname{WL}$) constructs a coloring $c=c_{m,{\mathscr{R}}}$ assigning a color $c(x)$ to each $m$-tuple $x\in\Omega^{m}$. The coloring is canonical in the sense that for any bijection $f:\Omega\to\Omega^{\prime}$, the induced bijection $\Omega^{m}\to{\Omega^{\prime}}{\phantom{x}{}^{m}}$ takes the coloring $c_{m,{\mathscr{R}}}$ to the coloring $c_{m,{\mathscr{R}}^{\prime}}$, where =${\mathscr{R}}^{\prime}=\{s^{f}:\ s\in{\mathscr{R}}\}$. More details (including the exact definition of the $m$-dim $\operatorname{WL}$) can be found in [15] (see also [17]).
@@
At this point, we assume that $m\geq 2$. For a ($2$-ary) coherent configuration ${\cal X}$, denote by $\operatorname{WL}_{m}({\cal X})$ the $m$-ary coherent configuration $\operatorname{WL}_{m}({\mathscr{R}})$ with ${\mathscr{R}}=S({\cal X})$. Following [28], we say that two coherent configurations ${\cal X}$ and ${\cal X}^{\prime}$ are $\operatorname{WL}_{m}$-equivalent with respect to an algebraic isomorphism $\varphi:{\cal X}\to{\cal X}^{\prime}$ if there exists an algebraic isomorphism $\hat{\varphi}:\operatorname{WL}_{m}({\cal X})\to\operatorname{WL}_{m}({\cal X}^{\prime})$ such that its $2$-projection $\hat{\varphi}_{2}$ extends $\varphi$, i.e., $\hat{\varphi}_{2}(s)=\varphi(s)$ for all $s\in S({\cal X})$. (Note that we always have $\operatorname{pr}_{2}\operatorname{WL}_{m}({\cal X})\geq{\cal X}$ and $\operatorname{pr}_{2}\operatorname{WL}_{m}({\cal X}^{\prime})\geq{\cal X}^{\prime}$.)
The $m$-dim $\operatorname{WL}$ algorithm identifies a coherent configuration ${\cal X}$ if for any algebraic isomorphism $\varphi:{\cal X}\to{\cal X}^{\prime}$ with respect to which ${\cal X}$ and ${\cal X}^{\prime}$ are $\operatorname{WL}_{m}$-equivalent, the set $\operatorname{Iso}({\cal X},{\cal X}^{\prime},\varphi)$ is not empty. The $\operatorname{WL}$-dimension $\dim_{\scriptscriptstyle\operatorname{WL}}({\cal X})$ of the coherent configuration ${\cal X}$ is defined to be the smallest positive integer $m\geq 2$ such that the $m$-dim $\operatorname{WL}$ algorithm identifies ${\cal X}$.
With any graph ${\cal G}$, we associate the coherent configuration $\operatorname{WL}({\cal G})=\operatorname{WL}_{2}({\mathscr{R}})$, where the set ${\mathscr{R}}$ is the singleton consisting of the edge set of ${\cal G}$. Let us define the $\operatorname{WL}$-dimension of ${\cal G}$ by the formula
(9) $\dim_{\scriptscriptstyle\operatorname{WL}}({\cal G})=\dim_{\scriptscriptstyle\operatorname{WL}}(\operatorname{WL}({\cal G})).$
@@ 4.2. Pebble games
It is convenient to express the $\operatorname{WL}_{m}$-equivalence of coherent configurations ${\cal X}$ and ${\cal X}^{\prime}$ with respect to an algebraic isomorphism $\varphi:{\cal X}\to{\cal X}^{\prime}$ in terms of the pebbling game ${\mathscr{C}}_{m}(\varphi)$, see [32, 6]. The idea and the main result of this subsection go back to a fundamental paper [4].
There are two players, called Spoiler and Duplicator, and $m$ pairwise distinct pebbles, each given in duplicate. The game consists of rounds and each round consists of two parts. At the first part, Spoiler chooses a set $A^{\prime}$ of points belonging to ${\cal X}$ or ${\cal X}^{\prime}$. Duplicator responds with a set $A$ in the other coherent configuration, such that $|A|=|A^{\prime}|$ (if this is impossible, then Duplicator loses). At the second part, Spoiler places one of the pebbles [^2] It is allowed to move previously placed pebbles to other vertices and place more than one pebble on the same vertex. on a point in $A$. Duplicator responds by placing the copy of the pebble on some point of $A^{\prime}$.
The configuration after a round is determined by a bijection $f:\Delta\to\Delta^{\prime}$, where $\Delta\subseteq\Omega$ (respectively, $\Delta^{\prime}\subseteq\Omega^{\prime}$) is the set of points in ${\cal X}$ (respectively, in ${\cal X}^{\prime}$), covered by pebbles, and any two points $\alpha\in\Delta$ and $\alpha^{f}\in\Delta^{\prime}$ are covered by the copies of the same pebble. Duplicator wins the round if
@@
The game starts from an initial configuration (which is considered as the configuration after zero rounds), i.e., a pair $(x,x^{\prime})\in\Omega^{k}\times{\Omega^{\prime}}{\phantom{x}{}^{k}}$ for which $k\leq m$ and the points $x_{i}$ and $x^{\prime}_{i}$ are covered with copies of the same pebble, $i=1,\ldots,k$. We say that Duplicator has a winning strategy for the game ${\mathscr{C}}_{m}(\varphi)$ on ${\cal X}$ and ${\cal X}^{\prime}$ with initial configuration $(x,x^{\prime})$ if, regardless of Spoiler’s actions, Duplicator wins after any number of rounds.
@@
Let ${\cal X}$ and ${\cal X}^{\prime}$ be coherent configurations on, respectively, $\Omega$ and $\Omega^{\prime}$, and let $\varphi:{\cal X}\to{\cal X}^{\prime}$ be an algebraic isomorphism. Assume that an integer $m\geq 2$ and a bijection $f:\Omega^{m}\to{\Omega^{\prime}}{\phantom{x}{}^{m}}$ are such that for each $x\in\Omega^{m}$, Duplicator has a winning strategy in the game ${\mathscr{C}}_{m+1}(\varphi)$ on ${\cal X}$ and ${\cal X}^{\prime}$ with initial configuration $(x,x^{\prime})$, where $x^{\prime}=x^{f}$. Then ${\cal X}$ and ${\cal X}^{\prime}$ are $\operatorname{WL}_{m}$-equivalent with respect to $\varphi$.
@@
Let $T\subseteq\Omega^{m}\times{\Omega^{\prime}}{\phantom{x}{}^{m}}$ consist of all pairs $(x,x^{\prime})$ such that Duplicator has a winning strategy in the pebble game ${\mathscr{C}}_{m+1}(\varphi)$ on ${\cal X}$ and ${\cal X}^{\prime}$ with initial configuration $(x,x^{\prime})$. By the assumption, $T$ contains each pair $(x,x^{f})$, $x\in\Omega^{m}$. Since $f$ is a bijection, the relation $T$ has full support. This completes the proof, because, by [32, Theorem 3.3], ${\cal X}$ and ${\cal X}^{\prime}$ are $\operatorname{WL}_{m}$-equivalent with respect to $\varphi$ if and only if $T$ has full support. ∎

### v2

@@ 3. Binary schemes
Let $\mathcal{X}$} be a scheme on $\Omega$. For any integer $m\geq 1$ and a tuple $x\in\Omega^{m}$, we define the $m\times m$ array
(7) $R_{{\mathcal{X}},m}(x)=(r_{\mathcal{X}}(x_{i},x_{j}):\ 1\leq i,j\leq m).$
The scheme $\mathcal{X}$} is said to be binary if for any $m$ and any two $m$-tuples $x,y\in\Omega^{m}$ such that $R_{{\mathcal{X}},m}(x)=R_{{\mathcal{X}},m}(y)$, there exists an automorphism of $\mathcal{X}$}, that takes $x$ to $y$. In this case, the group $\operatorname{Aut}({\mathcal{X}})$ is transitive, and acts transitively on each basis relation of $\mathcal{X}$}.
@@
Below, $\mathcal{X}$} is a scheme on $\Omega$, $m\geq 1$ is an integer, and $x,y\in\Omega^{m}$ are such that $R(x)=R(y)$, where $R=R_{{\mathcal{X}},m}$. First, assume that $\mathcal{X}$} is regular. Then there exists a unique automorphism $f\in\operatorname{Aut}({\mathcal{X}})$ such that ${x_{1}}^{f}=y_{1}$. Let $1\leq i\leq m$. By the assumption, the point $y_{i}$ is the unique one for which $r(x_{1},x_{i})=r(y_{1},y_{i})$, where $r=r_{\mathcal{X}}$. Since $r(x_{1},x_{i})^{f}=r(y_{1},x_{i}^{f})$, this shows that $y_{i}={x_{i}}^{f}$. Thus $y=x^{f}$ and $\mathcal{X}$} is binary.
Let ${\mathcal{X}}={\mathcal{X}}_{1}\otimes{\mathcal{X}}_{2}$, where ${\mathcal{X}}_{1}$ and ${\mathcal{X}}_{2}$ are binary schemes on $\Omega_{1}$ and $\Omega_{2}$, respectively. Denote by $\operatorname{pr}_{i}:\Omega_{1}\times\Omega_{2}\to\Omega_{i}$ the natural projection, $i=1,2$. Identifying in a natural way the sets $(\Omega_{1}\times\Omega_{2})^{m}$ and $\Omega_{1}^{\phantom{\,}m}\times\Omega_{2}^{\phantom{\,}m}$, we can write $x=(\operatorname{pr}_{1}x,\operatorname{pr}_{2}x)$ and $y=(\operatorname{pr}_{1}y,\operatorname{pr}_{2}y)$. By the definition of tensor product, we have
@@
where $R_{1}=R_{{\mathcal{X}}_{1},m}$, and $R_{2}=R_{{\mathcal{X}}_{2},m}$. Since the scheme ${\mathcal{X}}_{i}$ is binary, there exists an automorphism $f_{i}\in\operatorname{Aut}({\mathcal{X}}_{i})$ such that $\operatorname{pr}_{i}y=(\operatorname{pr}_{i}x)^{f_{i}}$, $i=1,2$. It follows that the permutation $f:=(f_{1},f_{2})$ of $\Omega_{1}\times\Omega_{2}$ belongs to the group $\operatorname{Aut}({\mathcal{X}}_{1})\times\operatorname{Aut}({\mathcal{X}}_{2})=\operatorname{Aut}({\mathcal{X}})$. Thus,
@@
and the scheme $\mathcal{X}$} is binary.
Now let ${\mathcal{X}}={\mathcal{X}}_{1}\wr{\mathcal{X}}_{2}$. Then $e={\bf 1}_{\Omega_{1}}\otimes 1_{\Omega_{2}}$ is a parabolic of the scheme $\mathcal{X}$}. Each class of $e$ is of the form $\bar{\alpha}=\Omega_{1}\times\{\alpha\}$, where $\alpha\in\Omega_{2}$. The bijection $\Omega_{2}\to\Omega/e$, $\alpha\mapsto\bar{\alpha}$, $\alpha\in\Omega_{2}$ is an isomorphism from the scheme ${\mathcal{X}}_{2}$ to the quotient scheme $\bar{\mathcal{X}}={\mathcal{X}}_{\Omega/e}$; in particular, the scheme $\bar{\mathcal{X}}$ is binary.
@@
where $\bar{R}=R_{\bar{\mathcal{X}},m}$, $\bar{x}=(\overline{x_{1}},\ldots,\overline{x_{m}})$, and $\bar{y}=(\overline{y_{1}},\ldots,\overline{y_{m}})$. Thus there exists an automorphism $\bar{f}\in\operatorname{Aut}(\bar{\mathcal{X}})$ such that $\bar{y}={\bar{x}}^{\bar{f}}$. Because the natural homomorphism $\operatorname{Aut}({\mathcal{X}})\to\operatorname{Aut}(\bar{\mathcal{X}})$ is surjective (see [5, Theorem 3.4.6.]), one can find an automorphism $g\in\operatorname{Aut}({\mathcal{X}})$ such that $g^{\Omega/e}=\bar{g}$. Replacing $x$ by $x^{g}$ and taking into account that $R(x^{g})=R(x)$, we may assume that
@@
where $R_{\Delta}=R_{{\mathcal{X}}_{\Delta},k}$. Because the scheme ${\mathcal{X}}_{\Delta}={\mathcal{X}}_{1}$ is binary, there exists an automorphism $f_{\Delta}\in\operatorname{Aut}({\mathcal{X}}_{\Delta})$ such that $y_{\Delta}=(x_{\Delta})^{f_{\Delta}}$. Let us define a permutation $f\in\operatorname{Sym}(\Omega)$ such that $f^{\Delta}=f_{\Delta}$ for all $\Delta\in\Omega/e$. Then obviously $x^{f}=y$. To complete the proof, it remains to note that $f\in\operatorname{Aut}({\mathcal{X}})$, because the kernel of the homomorphism $\operatorname{Aut}({\mathcal{X}})\to\operatorname{Aut}(\overline{{\mathcal{X}}})$ coincides with the direct product of $|\Omega/e|$ copies of the group $\operatorname{Aut}({\mathcal{X}}_{\Delta})$. ∎
@@ 4.1. Definitions
Let $\mathscr{R}$} be a linear ordered set of binary relations on $\Omega$, and let $m\geq 1$ be an integer. The Weisfeiler-Leman algorithm (the $m$-dim $\operatorname{WL}$) constructs a coloring $c=c_{m,{\mathscr{R}}}$ assigning a color $c(x)$ to each $m$-tuple $x\in\Omega^{m}$. The coloring is canonical in the sense that for any bijection $f:\Omega\to\Omega^{\prime}$, the induced bijection $\Omega^{m}\to{\Omega^{\prime}}{\phantom{x}\hskip-5.69054pt{}^{m}}$ takes the coloring $c_{m,{\mathscr{R}}}$ to the coloring $c_{m,{\mathscr{R}}^{\prime}}$, where ${\mathscr{R}}^{\prime}=\{s^{f}:\ s\in{\mathscr{R}}\}$. More details (including the exact definition of the $m$-dim $\operatorname{WL}$) can be found in [15] (see also [17]).
@@
At this point, we assume that $m\geq 2$. For a ($2$-ary) coherent configuration $\mathcal{X}$}, denote by $\operatorname{WL}_{m}({\mathcal{X}})$ the $m$-ary coherent configuration $\operatorname{WL}_{m}({\mathscr{R}})$ with ${\mathscr{R}}=S({\mathcal{X}})$. Following [28], we say that two coherent configurations $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ are $\operatorname{WL}_{m}$-equivalent with respect to an algebraic isomorphism $\varphi:{\mathcal{X}}\to{\mathcal{X}}^{\prime}$ if there exists an algebraic isomorphism $\hat{\varphi}:\operatorname{WL}_{m}({\mathcal{X}})\to\operatorname{WL}_{m}({\mathcal{X}}^{\prime})$ such that its $2$-projection $\hat{\varphi}_{2}$ extends $\varphi$, i.e., $\hat{\varphi}_{2}(s)=\varphi(s)$ for all $s\in S({\mathcal{X}})$. (Note that we always have $\operatorname{pr}_{2}\operatorname{WL}_{m}({\mathcal{X}})\geq{\mathcal{X}}$ and $\operatorname{pr}_{2}\operatorname{WL}_{m}({\mathcal{X}}^{\prime})\geq{\mathcal{X}}^{\prime}$.)
The $m$-dim $\operatorname{WL}$ algorithm identifies a coherent configuration $\mathcal{X}$} if for any algebraic isomorphism $\varphi:{\mathcal{X}}\to{\mathcal{X}}^{\prime}$ with respect to which $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ are $\operatorname{WL}_{m}$-equivalent, the set $\operatorname{Iso}({\mathcal{X}},{\mathcal{X}}^{\prime},\varphi)$ is not empty. The $\operatorname{WL}$-dimension $\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{X}})$ of the coherent configuration $\mathcal{X}$} is defined to be the smallest positive integer $m\geq 2$ such that the $m$-dim $\operatorname{WL}$ algorithm identifies $\mathcal{X}$}.
With any graph $\mathcal{G}$}, we associate the coherent configuration $\operatorname{WL}({\mathcal{G}})=\operatorname{WL}_{2}({\mathscr{R}})$, where the set $\mathscr{R}$} is the singleton consisting of the edge set of $\mathcal{G}$}. Let us define the $\operatorname{WL}$-dimension of $\mathcal{G}$} by the formula
(9) $\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{G}})=\dim_{\scriptscriptstyle\operatorname{WL}}(\operatorname{WL}({\mathcal{G}})).$
@@ 4.2. Pebble games
It is convenient to express the $\operatorname{WL}_{m}$-equivalence of coherent configurations $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ with respect to an algebraic isomorphism $\varphi:{\mathcal{X}}\to{\mathcal{X}}^{\prime}$ in terms of the pebbling game ${\mathscr{C}}_{m}(\varphi)$, see [32, 6]. The idea and the main result of this subsection go back to a fundamental paper [4].
There are two players, called Spoiler and Duplicator, and $m$ pairwise distinct pebbles, each given in duplicate. The game consists of rounds and each round consists of two parts. At the first part, Spoiler chooses a set $A^{\prime}$ of points belonging to $\mathcal{X}$} or $\mathcal{X}$$^{\prime}$. Duplicator responds with a set $A$ in the other coherent configuration, such that $|A|=|A^{\prime}|$ (if this is impossible, then Duplicator loses). At the second part, Spoiler places one of the pebbles [^2] It is allowed to move previously placed pebbles to other vertices and place more than one pebble on the same vertex. on a point in $A$. Duplicator responds by placing the copy of the pebble on some point of $A^{\prime}$.
The configuration after a round is determined by a bijection $f:\Delta\to\Delta^{\prime}$, where $\Delta\subseteq\Omega$ (respectively, $\Delta^{\prime}\subseteq\Omega^{\prime}$) is the set of points in $\mathcal{X}$} (respectively, in $\mathcal{X}$$^{\prime}$), covered by pebbles, and any two points $\alpha\in\Delta$ and $\alpha^{f}\in\Delta^{\prime}$ are covered by the copies of the same pebble. Duplicator wins the round if
@@
The game starts from an initial configuration (which is considered as the configuration after zero rounds), i.e., a pair $(x,x^{\prime})\in\Omega^{k}\times{\Omega^{\prime}}{\phantom{x}\hskip-5.69054pt{}^{k}}$ for which $k\leq m$ and the points $x_{i}$ and $x^{\prime}_{i}$ are covered with copies of the same pebble, $i=1,\ldots,k$. We say that Duplicator has a winning strategy for the game ${\mathscr{C}}_{m}(\varphi)$ on $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ with initial configuration $(x,x^{\prime})$ if, regardless of Spoiler’s actions, Duplicator wins after any number of rounds.
@@
Let $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ be coherent configurations on, respectively, $\Omega$ and $\Omega^{\prime}$, and let $\varphi:{\mathcal{X}}\to{\mathcal{X}}^{\prime}$ be an algebraic isomorphism. Assume that an integer $m\geq 2$ and a bijection $f:\Omega^{m}\to{\Omega^{\prime}}{\phantom{x}\hskip-5.69054pt{}^{m}}$ are such that for each $x\in\Omega^{m}$, Duplicator has a winning strategy in the game ${\mathscr{C}}_{m+1}(\varphi)$ on $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ with initial configuration $(x,x^{\prime})$, where $x^{\prime}=x^{f}$. Then $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ are $\operatorname{WL}_{m}$-equivalent with respect to $\varphi$.
@@
Let $T\subseteq\Omega^{m}\times{\Omega^{\prime}}{\phantom{x}\hskip-5.69054pt{}^{m}}$ consist of all pairs $(x,x^{\prime})$ such that Duplicator has a winning strategy in the pebble game ${\mathscr{C}}_{m+1}(\varphi)$ on $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ with initial configuration $(x,x^{\prime})$. By the assumption, $T$ contains each pair $(x,x^{f})$, $x\in\Omega^{m}$. Since $f$ is a bijection, the relation $T$ has full support. This completes the proof, because, by [32, Theorem 3.3], $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ are $\operatorname{WL}_{m}$-equivalent with respect to $\varphi$ if and only if $T$ has full support. ∎